OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x +3y = 4.

OpenStudy (saifoo.khan):

-4 = -2(1)/3 + c c = -10/3

OpenStudy (saifoo.khan):

y = mx + c y = -2x/3 -10/3 3y = -2x -10